OpenStudy (anonymous):

I REALLY NEED HELP, FORGOT HOW TO DO SIMPLE PROBLEM #2 Solve the linear equations. Identify the equation as being conditional, the indentity, or a contradiction( I do know how to tell which one of those is which) (3x+1)/2 - 4/5 = x/10 + (2x-3)/5

OpenStudy (mathmale):

Hi, J-Man, I'd suggest simplifying our situation by identifying the LCD of those several fractions within your equation and then multiplying every term by that LCD. This will eliminate fractions and make it easier to simplify the equation further.

OpenStudy (anonymous):

That's where I have my problem I forgot what to do exactly when I have the LCD

OpenStudy (mathmale):

What LCD have you come up with? Here's a rough idea of what those descriptors mean: "conditional": The equation is true only for a certain value of x. "identity": always true, regardless of the value of x.

OpenStudy (mathmale):

"contradition": never true, for example, 0x + 0y = 7

OpenStudy (anonymous):

I just forgot how to find the LCD and I don't know why. >.< I tried multiplying both of the sides by 2 and by 5 and I guess I don't know what I am doing anymore

OpenStudy (mathmale):

Your equation has four terms. All but one of these terms have denominators. Rather than go into a lot of detail, I'll just mention that both 2 and 5 divide into the largest denominator, 10. Thus, the LCD is 10. Multiply every term of the equation b y the LCD: 10

OpenStudy (mathmale):

\[10(\frac{ 3x+1 }{ 2 }-\frac{ 4 }{ 5 }=\frac{ x }{ 10 }+\frac{ 2x-3 }{ 5 })\] Simplify everywhere possible. for example, in the first term, 10/2 = 5, so we have 5(3x+1). What about the other 3 terms? After simplification in this manner, try solving for x. Is this possible?

OpenStudy (anonymous):

Oh okay for some reason I've been thinking Lowest Common Den instead of Largest. this is another part where I am confused for some reason. would the first part be like: (30x+10)/20 or what. I wish I knew why I can't remember this stuff

OpenStudy (anonymous):

Oh I see, I've been doing it completely wrong. How did you get that

OpenStudy (mathmale):

I assume you're speaking of the first TERM of your equation after multiplying everything by 10. \[10\frac{ 3x+1 }{ 2 }=5(3x+1)=15x+5.\]

OpenStudy (mathmale):

It's important you work through this yourself. But the next 3 terms are, in order, -8, x and 4x-6. combine all four terms (that is, combine like terms). Then try solving for x.

OpenStudy (anonymous):

I guess I'm just not too good with this. What happened to the -4/5

OpenStudy (anonymous):

I will be right back if you are still on then

OpenStudy (mathmale):

Probably won't be, as I have errands to do; but I hope to work with you again!

OpenStudy (anonymous):

No worries Thank you very much the help and getting me started

OpenStudy (anonymous):

@mathmale after reading back through your replies on this question I finally understood again! Thank you so much!

OpenStudy (mathmale):

I'm so happy to hear that! Thanks for taking the trouble to let me know your status! Keep up the good work. Hope to "see" you again online soon.